OpenStudy (anonymous):

Do the equations 3+y=20 and x-3y=7 represent two perpendicular lines? Why or Why not?

jimthompson5910 (jim_thompson5910):

If it's 3x+y=20 and x-3y=7, then they are two perpendicular lines

OpenStudy (anonymous):

but how?

jimthompson5910 (jim_thompson5910):

Solve for y in each case 3x+y=20 y=-3x+20 Slope here is -3 ----------------- x-3y=7 -3y=-x+7 y = 1/3x-7/3 Slope here is 1/3 Since the two slopes are negative reciprocals of one another, the two lines are perpendicular to each other.